# Empirical analysis of Binding Precedent efficiency in the Brazilian Supreme Court via Similar Case Retrieval

### v2

@@ 4.5 BP 26
The trends of the TF-IDF+SVM or TF-IDF+logistic curves, in Fig. 15, can be mapped to exact juridical events. First, the peak around 02/23/2006 corresponds to the declaration of unconstitutionality in HC 82.959/SP, the case that launched the discussion leading up to BP 26. The almost immediate reaction of the Brazilian Congress to address the unconstitutionality declared by this judgment, with the presentation of a law project as early as 03/23/2006, would explain the subsequent decline.
@@
Additional metadata for documents predicted by TF-IDF+SVM for BP 26 Refer to caption
@@
Figure 18: First: Type of legal process used in the documents predicted by the model TF-IDF+SVM for BP 26, among RE (Recurso Extraordinário), Rcl (Reclamação) and HC (Habeas Corpus). Second: State to which these cases belong, with label “UNKNOWN” when the metadata is unavailable. Third: Type of decision taken at trial: processual if deferred for procedural matters, or deferido (accepted) and indeferido (rejected). Fourth: Presence of the words exame criminológico (criminal examination) and fundamentado (grounded/founded).
@@
Fig. 18 reveals other important information: starting from 2017, the cases tend to be deferred by the High Court based on procedural matters. More precisely, we have observed that it is common for cases to arrive in front of the judges having already lost their purpose (for instance because the convicted, in the meantime, obtained regime progression through other means). But more frequently cases are deferred because the STF considered that reclamação was not the correct procedural instrument. Even though the chance of an appeal being accepted might be low, we could reasonably think that defenders continue submitting them, gambling on this chance. This constitutes a typical case of the New avenue to the court and procedural issues hypothesis.
@@
In conclusion, our analysis shows that the expected effect of unifying the jurisprudence surrounding BP 26 — regime progression for convicted of heinous crimes — was not entirely achieved. The possibility of the progression is taken for granted, but it has simultaneously provoked a new wave of discussions, centered around the use of criminological examination, and more precisely the justification of its request. In this respect, the vague wording of BP 26 is to blame. However, we should note that vague wording is a common trait of legal provisions. Under certain conditions, it may be viewed not as a problem inherent to the production of Law, but as a quality of it. Vagueness in legal text allows for flexibility of interpretation and may be intentionally used when relevant context-specific dimensions of a legal problem are not known ex ante. Thus, when a general or vague term is used in Law, aspects of the solution of the problem are, in practice, deferred to be solved at a posterior time, when more information about the concrete situation is available to the parties involved. Vagueness in the legal text is both unavoidable, to a certain extent, and desirable, as precision can generate arbitrary solutions that are contradictory to the main purpose of a norm under specific or unforeseen circumstances. As Endicott [22, p. 14] puts it: “Far from being repugnant to the idea of making a norm, vagueness is of central importance to law-makers (…) It is a central technique of normative texts.”
@@
When elaborating a binding precedent, STF is in the unique position of being both the creator of a general normative text and the ultimate interpreter of that same text. This induces legal tensions that our study of BP 26 has brought to light: if the vague wording of BP 26 has allowed for more flexibility in the determination of possible uses of criminological examination, it has also caused many cases to be brought to STF for clarification, rendering one of the main purposes of a BP — that of reducing legal disputes and legal uncertainty — ineffective.
@@ 4.6 BP 37
Many legal cases awaiting judgment by the Justices of STF have been recognized as having general repercussion. In August 2014, the Justices ruled on one of these cases (the extraordinary appeal RE 592.317), which would release 1100 ordinary cases for judgment in lower judiciary branches. In this appeal, the STF revisited and applied one of its precedents to overrule a decision from a Brazilian State court. This (non-binding) precedent, Súmula 339 (Precedent 339), establishes that the judiciary cannot increase the salaries of public servants on the grounds of isonomy. During the debates on the mentioned appeal, Justice Gilmar Mendes suggested converting Precedent 339 into a binding precedent, given its consistent application in STF’s decisions. This way, lower judiciary branches would have to rule according to what was established by the Supreme Court, and hopefully, that would reduce the number of cases related to it that reach STF. In October 2014, BP 37 was created with the same wording as Precedent 339:
@@
Long before the creation of this binding precedent, a law from 2003 (Law 10,698) was enacted to grant an increase of R\$ 59.87 to all federal public servants. This resulted in 30,000 legal cases over 15 years arguing that the adjustment should not be a fixed amount but rather proportional to each servant’s salary (specifically 13.23%) [^31] https://www.conjur.com.br/2017-out-23/lei-deu-aumento-59-servidores-produziu-30-mil-processos. Different judges issued different rulings, so legal uncertainty took place. Many cases reached the STF, and the Justices, opposed to the adjustment, ended up citing Precedent 339 and later Binding Precedent 37 in their decisions. While BP 37 and repeated rulings from the STF demonstrated the unconstitutionality of such an adjustment, the controversy surrounding the issue continued to create some degree of legal uncertainty. As a result, the number of decisions citing this BP throughout time remained high.
Mathematical results
We show in Fig. 19 the predictions of the five models, as well as a regular expression search for the words isonomia, vencimentos and servidores públicos (named “regex” in the legend), and a search for Súmula 339 (Precedent 339). For BP 37, we tuned the models’ thresholds as those such that at least 80% are predicted, and not 90%, as is the case for BPs 11, 14 and 26 (see LABEL:tab:adjusted_probabilities). In fact, BP 37 is the precedent that has the largest number of documents citing it; choosing this smaller threshold means that fewer documents are detected, and we conducted a better analysis with this value. As demonstrated in the bottom part of the figure, the TF-IDF models presented the best results, along with regex, since they faithfully follow the groundtruth curve after the BP’s publication.
Models’ predictions for BP 37 Refer to caption
Figure 19: Number of documents predicted by each model for BP 37 in Dataset #2, represented as a histogram (window length of six months) and interpolated via quadratic spline. We give the predictions for the thresholds adapted to Dataset #1 (dashed) and #2 (solid), as described in LABEL:tab:adjusted_probabilities. Two views are given, the second one zooming in on the ordinate axis. In addition, we represent a regex query for Súmula 339.
One observes a few common trends in the figure. Before the publication of the BP, the model LSTM, as well as the regex search, shows a clear peak in 2005. These documents indeed contain words that characterize BP 37 and related decisions. A manual inspection shows that these documents involve Law 8,622/1993, granting an increase of 28.86% in the salary of some categories of public servants. The documents are essentially appeals, raised by servants who, although belonging to categories not foreseen by this law, were requesting the same increase. As these servants were not subject to Precedent 339, the precedent did not apply, and the High Court granted them a salary increase. In summary, these predicted documents seem somehow related to the BP’s topic and cases that cite it, but do not refer to the application of Precedent 339. Ideally, they should not have been detected. The TF-IDF models did not predict them, reinforcing their good performance in the task. We point out that all the documents predicted by the TF-IDF models in 2004 actually mention Precedent 339. As this precedent and the BP have the same wording, these documents were correctly predicted.
Besides, after the publication of BP 37, two peaks in the number of predictions can be observed (in 2016 and 2017), reflected by all the models. As it turns out, these peaks can be mapped to specific juridical events, as we will study further in the juridical discussion below (see Fig. 21).
In terms of quantity, LSTM is the model with the highest number of predictions before the BP publication date, switching places with BERT afterward. Reading a sample of these documents shows that LSTM and BERT’s predictions are only vaguely related to the theme of BP 37. Curiously, the TF-IDF models predicted almost no document before 2012. One could say that TF-IDF is overfitted (that is, only estimate documents citing the BP), while the other models are underfitted (that is, estimate non-relevant documents). To analyze the reasons for this discrepancy, we give in Fig. 20 the correlations between a few relevant words in the BP’s wording: aumentar, cabe, isonomia, legislativa, públicos, servidores and vencimentos (to increase, fits/suits, isonomy, legislative, public, servants, salaries). As expected, the documents citing BP 37 show a significant correlation between servidores and públicos (as in “public servants”, a common expression in Portuguese), but also, more interestingly, aumentar and legislativa. As observed in the figure, all the models were able to catch these correlations. However, presenting these correlations is not enough for a model to give correct predictions: for instance, only 47.6% of LSTM’s predictions contain the expression “public servants”, compared to 78.7%, 83.5% and 86.9% for the TF-IDF models. The large number of predictions from LSTM and BERT after 2014, following an increasing trend, simply suggests a poor generalization of these models — we recall that when validated on Dataset #1, they obtained good results (see LABEL:tab:similar_case_matching).
Correlations between words among the models’ predictions for BP 37 Refer to caption
Figure 20: Correlations between the selected relevant words, when restricted to documents citing BP 37 (“Groundtruth”) or documents predicted by one of our models.
To understand the low number of predictions of TF-IDF before 2014, compared to the documents citing Precedent 339, one can study the important words of TF-IDF, collected in LABEL:tab:importance_features. In particular, isonomia has the greatest importance, hence we expect that a large portion of the predicted documents contain this word. It is indeed the case, at a rate of 79.1%, 83.6%, and 88.9% for SVM, logistic, and randomforest. Comparatively, only 63.9% of the documents citing Precedent 339 mention this word. This suggests that the model is biased towards estimating principally documents employing isonomia, which is not representative of all the documents similar to BP 37, explaining the small number of documents predicted.
Juridical discussion
@@
As a consequence of the above analysis, the predictions of TF-IDF+randomforest are only trustworthy after the publication of BP 37. Fortunately, similar documents before the publication can be studied through citations of Precedent 339. We will therefore divide our discussion into two parts, first studying the predictions of TF-IDF+randomforest. We note that it would have been equivalent, as far as post-publication documents are concerned, to consider the groundtruth curve (actual citations of BP 37, see Fig. 19).
@@
Fig. 21 gathers relevant metadata regarding TF-IDF+randomforest’s predictions: the type of process, the state of emission, and the mention of the words 1987, 10.698, Mogi Guaçu and magistratura, which we will explain throughout the text. The first peak of the curve, in 2016, can be related to real-world juridical events as follows. As depicted in the figure, most predicted documents in that year were Extraordinary Appeals (RE) that came from the State of Rio de Janeiro (RJ). A manual inspection of some of these documents shows the application of BP 37 in the context of the RJ State Law 1,206/1987. This is reflected by the citation curve for “1987”. Given that this law was enacted to increase the salaries of public servants from the Executive and Legislative branches, servants from the Judiciary questioned their exclusion by taking legal action. After citing BP 37 in several decisions, STF issued a final ruling dismissing the servants’ request in October 2016323232General repercussion thesis ARE 909,437 https://redir.stf.jus.br/paginadorpub/paginador.jsp?docTP=TP&docID=11828219.
@@
Figure 21: Top: State to which the legal process used in the documents predicted by TF-IDF+randomforest for BP 37 belong, among the top three. Bottom: Documents containing the words 1987, 10.698, Mogi Guaçu or magistratura.
The second peak of decisions, from 2017 onwards, can be understood as a combination of several initiatives. The first reason is Law 13,317/2016 which altered the career and salary structure of public servants. This led to a significant increase in cases claiming a 13.23% salary adjustment, based on Law 10.698/2003, as shown by the citation curve in Fig. 21 (see citations of “10.698”). As a response, Justice Gilmar Mendes proposed in 2017 the creation of a new and specific binding precedent that would explicitly prohibit the granting of the mentioned adjustment [^33] https://www.conjur.com.br/dl/psv-128-reajuste-1323.pdf. After further analysis of the proposition, which considered the Prosecutor General’s manifestation [^34] https://portal.stf.jus.br/processos/downloadPeca.asp?id=312747008&ext=.pdf and others, the creation of this new BP was rejected.
In addition, the many mentions of “Mogi Guaçu”, a municipality in São Paulo, can be attributed to a collective agreement signed between the municipality and workers’ syndicate (Lei Complementar Municipal 1.121/2011353535https://sistema.camaramogiguacu.sp.gov.br/consultas/norma_juridica/norma_juridica_mostrar_proc?cod_norma=5779), determining the incorporation of bonuses into the salaries of all municipal employees and civil servants. A somewhat similar movement can be observed through the documents citing magistratura (members of the judiciary, such as judges), which relates to a controversy concerning the licença prêmio, a bonus awarded every five years to public servants (federal, municipal or state). Certain members of the judiciary invoked the principle of isonomy to obtain this bonus, which was rejected, according to the terms of BP 37363636https://portal.stf.jus.br/jurisprudenciaRepercussao/verPronunciamento.asp?pronunciamento=7188706.
In conclusion, the four trends observed on the citations of BP 37 — salary increases for public servants from Law 1,206/1987 or Law 13,317/2016, the Mogi-Guaçu syndicate and licença prêmio for judiciary — are protests led by specific groups, as formulated by our hypothesis Resistance by a group of litigants or regional specificity.
@@
Figure 22: Top: Type of legal process used in the documents citing Precedent 339, among RE (Recurso Extraordinário), ARE (Recurso Extraordinário com Agravo) and AI (Agravo de Instrumento). Bottom: Search for the presence of the words Santa Catarina (the state), 10.192, emenda consitucional (constitutional amendment) or revisão geral anual (annual general revision).
@@
A first trend can be observed between 1997 and 2002: most of the documents mention the state “Santa Catarina”, as visualized on the blue curve. A manual inspection revealed that most documents refer to two general cases that discuss permanent increases in the salaries of state servants of Santa Catarina who have temporarily held commissioned positions (Complementary State Law 43/1992 and State Law 9,847/1995). Thus, citing Precedent 339 in its arguments, the State of Santa Catarina filed extraordinary appeals to the STF against lower court decisions that guaranteed to the servants (i) the right to the permanent increase or (ii) equal salary adjustment to those servants receiving such an increase due to past commissioned positions and those who are currently in equivalent positions.
@@
To continue, one observes a significant rise in citations between 2003 and 2010, made of three peaks. They are all linked to specific event, represented by the presence of the words 10.192, emenda consitucional (constitutional amendment) or revisão geral anual (annual general review). Regarding the first one, a manual inspection revealed they refer to the application of Precedent 339 in the context of Law 10,192/2001, a law that increased the salaries of non-public servants (10,87%) and which was used by public servants to require a similar increase.
@@
As far as the 2005 peak is concerned, it must be traced back to 1993. Indeed, that was the year the Laws 8,622 (already mentioned) and 8,627 came into effect, adjusting the salaries of higher ranks in the armed forces by 28.86%. STF Justices then recognized the same right to lower-ranking military personnel, understanding it was a case of general revision of salaries, therefore not applying Precedent 339 but rather Article 37, Section X of the Constitution. The Union filed extraordinary appeals373737such as AI 551.744 https://portal.stf.jus.br/processos/detalhe.asp?incidente=2308620 (Recursos extraordinários), and one of the arguments was that the Constitution allowed for different adjustments for different ranks (Article 142), but the Justices upheld their decision based on Article 37. Article 37 was amended by Constitutional Amendment 19/1998383838https://www.planalto.gov.br/ccivil_03/constituicao/Emendas/Emc/emc19.htm#art3, which established the necessity of a specific law for any form of salary setting or change, thus exempting the requirement for adjustments to consistently follow the same percentage. In light of this, the Federal Government requested, and the STF accepted, that the 28.86% adjustment should be valid only until June 1998, when the Constitutional Amendment came into effect.
@@
We now turn to the last peak. In addition to the need for a specific law for any form of setting or change in the salary of public servants, Constitutional Amendment 19/1998 also established that the general revision of salaries must occur annually and based on annual laws proposed by the Head of the Executive branch (President of the Republic or state governors). A manual inspection of documents between 2006 and 2008 found extraordinary appeals393939such as RE 524.145 https://portal.stf.jus.br/processos/detalhe.asp?incidente=2461438 requesting compensation for public servants due to omissions in issuing such a law. STF denied these appeals based on Precedent 339. After many years of judgments on this matter, STF stated in 2019 that not submitting a law’s proposition regarding the annual revision of public servants’ salaries does not create a subjective right to compensation; however, the Executive branch must provide justification for why the revision was not proposed (STF Tema 19404040https://portal.stf.jus.br/jurisprudenciaRepercussao/tema.asp?num=19[56].
@@
It has also created, due to the persistence of this resistance, an influx of cases to the higher courts, not only to STF, but also to STJ (Superior Tribunal de Justiça), corresponding to the hypothesis New avenue to the court. As the controversy around this issue ensues, new cases continue to be brought before STF, as new attempts to adjust the salary of public servants occur in the absence of formal laws. The contradiction of this never-ending administrative practice to the content of the BP, in itself, creates a clear avenue for the issue to be brought up time and time again before the court. In this case, however, it is interesting to note that this new avenue is not a consequence of lawyers trying to bend formal restrictions or take advantage of indirect legal connections to gain access to the court. Rather, this avenue is a result of direct confrontation between persevering administrative practices and what has been determined as the correct legal interpretation of the Law by STF.

### v3

@@ Discussion
The trends of the prediction curves, in Fig. 14, can be mapped to exact juridical events. First, the peak around 02/23/2006 corresponds to the declaration of unconstitutionality in HC 82.959/SP, the case that launched the discussion leading up to BP 26. The almost immediate reaction of the Brazilian Congress to address the unconstitutionality declared by this judgment, with the presentation of a law project as early as 03/23/2006, would explain the subsequent decline.
@@
Additional metadata for documents predicted by Longformer for BP 26 Refer to caption
@@
Figure 21: First: Type of legal process used in the documents predicted by the model Longformer for BP 26, among AI (Agravo de Instrumento), Rcl (Reclamação), and HC (Habeas Corpus). Second: State to which these cases belong, with label “UNKNOWN” when the metadata is unavailable. Third: Type of decision taken at trial: processual if deferred for procedural matters, or deferido (accepted) and indeferido (rejected). Fourth: Presence of the words exame criminológico (criminal examination) and fundamentado (grounded/founded). In all graphs, the bins have a length of six months. The solid vertical line represents the date of publication of BP 26, and the dashed one the trial of HC 82.959/SP (declaration of inconstitutionality).
@@
Fig. 21 reveals other important information: starting from 2017, the cases tend to be deferred by the Supreme Court based on procedural matters. More precisely, we have observed that it is common for cases to arrive in front of the judges having already lost their purpose (for instance because the convicted, in the meantime, obtained regime progression through other means). But more frequently, cases are deferred because the court considered that reclamação was not the correct procedural instrument. Even though the chance of an appeal being accepted might be low, we could reasonably think that defenders continue submitting them, gambling on this chance. This constitutes a typical case of the New avenue to the court and procedural issues hypothesis.
@@
In conclusion, our analysis shows that the expected effect of unifying the jurisprudence surrounding BP 26 — regime progression for convicted of heinous crimes — was not entirely achieved. The possibility of the progression is taken for granted, but it has simultaneously provoked a new wave of discussions, centered around the use of criminological examination, and more precisely the justification of its request. In this respect, the vague wording of BP 26 is to blame. However, we should note that vague wording is a common trait of legal provisions. Under certain conditions, it may be viewed not as a problem inherent to the production of Law, but as a quality of it. Vagueness in legal text allows for flexibility of interpretation and may be intentionally used when relevant context-specific dimensions of a legal problem are not known ex ante. Thus, when a general or vague term is used in Law, aspects of the solution of the problem are, in practice, deferred to be solved at a posterior time, when more information about the concrete situation is available to the parties involved. Vagueness in the legal text is both unavoidable, to a certain extent, and desirable, as precision can generate arbitrary solutions that are contradictory to the main purpose of a norm under specific or unforeseen circumstances. As Endicott [23, p. 14] puts it: “Far from being repugnant to the idea of making a norm, vagueness is of central importance to law-makers (…) It is a central technique of normative texts.”
@@
When elaborating a binding precedent, the STF is in the unique position of being both the creator of a general normative text and the ultimate interpreter of that same text. This induces legal tensions that our study of BP 26 has brought to light: if the vague wording of the precedent has allowed for more flexibility in the determination of possible uses of criminological examination, it has also caused many cases to be brought to the STF for clarification, rendering one of the main purposes of a BP — that of reducing legal disputes and legal uncertainty — ineffective.
@@ 5.6 BP 37
Many legal cases awaiting judgment by the Justices of STF have been recognized as having general repercussions. In August 2014, the Justices ruled on one of these cases (the Extraordinary Appeal RE 592.317383838RE 592.317 https://redir.stf.jus.br/paginadorpub/paginador.jsp?docTP=TP&docID=7181942), which would release 1100 ordinary cases for judgment in lower judiciary branches. In this appeal, the STF revisited and applied one of its precedents to overrule a decision from a Brazilian State court. This (non-binding) precedent, Súmula 339393939Precedent 339 https://portal.stf.jus.br/jurisprudencia/sumariosumulas.asp?base=30&sumula=1484 (Precedent 339), establishes that the judiciary cannot increase the salaries of public servants on the grounds of isonomy. During the debates on the mentioned appeal, Justice Gilmar Mendes suggested converting the Precedent into a binding precedent, given its consistent application in the STF’s decisions. This way, lower judiciary branches would have to rule according to what was established by the Supreme Court, which, hopefully, would reduce the number of cases related to it that reach the STF. In October 2014, BP 37 was created with the same wording as Precedent 339:
@@
Long before the creation of this Binding Precedent, a law from 2003 (Law 10.698404040Law 10.698/2003 https://www.planalto.gov.br/ccivil_03/leis/2003/l10.698.htm) was enacted to grant an increase of R\$ 59.87 to all federal public servants. This resulted in 30,000 legal cases over 15 years arguing that the adjustment should not be a fixed amount but rather proportional to each servant’s salary [^41] Report on Law 10.698 https://www.conjur.com.br/2017-out-23/lei-deu-aumento-59-servidores-produziu-30-mil-processos (specifically, 13.23%). Different judges issued different rulings, so legal uncertainty took place. Many cases reached the STF, and the Justices, opposed to the adjustment, ended up citing Precedent 339 and later Binding Precedent 37 in their decisions. While BP 37 and repeated rulings from the STF demonstrated the unconstitutionality of such an adjustment, the controversy surrounding the issue continued to create some degree of legal uncertainty. As a result, the number of decisions citing this BP throughout time remained high.
Discussion
@@
As a consequence of the analysis of the predicted documents in Section 4.2 (see Fig. 15), the models detect documents of interest pre- and post-publication of BP 37. In addition, similar documents before its publication can be studied through citations of Precedent 339. We will therefore divide our discussion into two parts, first studying the predictions of TF-IDF+random forest. We note that it would have been equivalent, as far as post-publication documents are concerned, to consider the groundtruth curve (actual citations of BP 37), or that of regex.
Fig. 22 gathers relevant metadata regarding TF-IDF+random forest’s predictions: the type of process, the state of emission, and the mention of particular words, which we will explain throughout the text. The first peak of the curve, in 2015, concerns complaints/appeals (Rcl) originating from the State of Acre (Fig. 22 (Top, Middle)). A manual inspection of some of these documents shows the application of BP 37 in the context of the Acre Complementary State Laws 58/1998424242Acre Complementary State Law 58/1998 https://www.al.ac.leg.br/leis/?p=11843 and 67/1999434343Acre Complementary State Law 67/1999 https://www.al.ac.leg.br/leis/?p=11843 (see the respective curves in the figure). While temporary teachers from that State are subject to the first law, the latter applies to tenured professionals. The main reason these two laws and BP 37 appear together in many documents is that a lower court in Acre granted 45 days of vacation to temporary teachers (the same duration granted to tenured teachers), basing its decision not on the laws but on the principle of isonomy. The State subsequently filed complaints to the STF to overturn the Court’s decision, arguing that BP 37 had not been respected. However, the STF rejected the complaints, stating that the BP pertains to salary increases, which was not relevant in this case.
@@
The second peak of the curve, in 2016, can be related to real-world juridical events as follows. As depicted in the figure, most predicted documents in that year were Extraordinary Appeals (RE) that came from the State of Rio de Janeiro (RJ). A manual inspection of some of these documents shows the application of BP 37 in the context of the RJ State Law 1.206/1987444444RJ State Law 1.206/1987 https://leisestaduais.com.br/rj/lei-ordinaria-n-1206-1987-rio-de-janeiro-dispoe-sobre-o-reajuste-de-vencimentos-e-proventos-do-funcionalismo-estadual-e-da-outras-providencias?origin=instituicao. This is reflected by the citation curve for “1.206”. Given that this law was enacted to increase the salaries of public servants from the Executive and Legislative branches, servants from the Judiciary questioned their exclusion by taking legal action. After citing BP 37 in several decisions, the STF issued a final ruling dismissing the servants’ request in October 2016454545General repercussion thesis ARE 909.437 https://redir.stf.jus.br/paginadorpub/paginador.jsp?docTP=TP&docID=11828219.
@@
Figure 22: Top: Type of legal process used in the documents predicted by the model TF-IDF+random forest for BP 37, among Rcl (Reclamação), RE (Recurso Extraordinário), and ARE (Recurso Extraordinário com Agravo). Middle: State to which these cases belong, with label “UNKNOWN” when the metadata is unavailable. Bottom: Documents containing the words 58/1998, 67/1999, 1.206, 10.698, Mogi Guaçu, or magistratura (Judiciary members). The bins have a length of six months.
The third peak of decisions, from 2017 onwards, can be understood as a combination of several initiatives. The first reason is Law 13.317/2016464646Law 13.317/2016 https://www.planalto.gov.br/ccivil_03/_ato2015-2018/2016/lei/l13317.htm which altered the career and salary structure of public servants. This led to a significant increase in cases claiming a 13.23% salary adjustment, based on Law 10.698/2003 (already mentioned), as shown by the citation curve in Fig. 22 (see citations of “10.698”). As a response, Justice Gilmar Mendes proposed in 2017 the creation of a new and specific binding precedent that would explicitly prohibit the granting of the mentioned adjustment [^47] Proposal for a BP by Justice Gilmar Mendes https://www.conjur.com.br/dl/psv-128-reajuste-1323.pdf. After further analysis of the proposition, which considered the Prosecutor General’s manifestation [^48] Proposal for BP 128 https://portal.stf.jus.br/processos/downloadPeca.asp?id=312747008&ext=.pdf and others, the creation of this new BP was rejected.
In addition, the many mentions of “Mogi Guaçu”, a municipality in São Paulo, can be attributed to a collective agreement signed between the municipality and workers’ syndicate (Lei Complementar Municipal 1.121/2011494949Mogi Guaçu Complementary State Law 1.121/2011 https://sistema.camaramogiguacu.sp.gov.br/consultas/norma_juridica/norma_juridica_mostrar_proc?cod_norma=5779), determining the incorporation of bonuses into the salaries of all municipal employees and civil servants. A somewhat similar movement can be observed through the documents citing magistratura (members of the judiciary, such as judges), which relates to a controversy concerning the licença prêmio (long-service leave), a bonus awarded every five years to public servants (federal, municipal, or state). Certain members of the judiciary invoked the principle of isonomy to obtain this bonus, which was rejected, in the terms of BP 37505050Statement on general repercussion, Justice Luiz Fux, the 2017-07-26 https://portal.stf.jus.br/jurisprudenciaRepercussao/verPronunciamento.asp?pronunciamento=7188706.
In conclusion, the five trends observed on the citations of BP 37 — vacation for temporary teachers in the State of Acre, salary increases for public servants from Law 1.206/1987 or Law 13.317/2016 in the State of Rio de Janeiro, the Mogi-Guaçu syndicate, and licença prêmio for judiciary — are protests led by specific groups, as formulated by our hypothesis Resistance by a group of litigants or regional specificity.
@@
Figure 23: Top: Type of legal process used in the documents citing Precedent 339, among Rcl (Reclamação), RE (Recurso Extraordinário), and ARE (Recurso Extraordinário com Agravo). Middle: State to which these cases belong. Bottom: Search for the presence of the words 10.192, emenda consitucional (constitutional amendment), militares (military personnel), or revisão geral anual (annual general revision). In all three graphs, the bins are six months long.
@@
A first trend can be observed between 1997 and 2002: most of the documents come from the State of Santa Catarina, as visualized on the blue curve in the middle graph. A manual inspection revealed that most documents refer to two general cases that discuss permanent increases in the salaries of state servants of Santa Catarina who have temporarily held commissioned positions (Complementary State Law 43/1992 and State Law 9.847/1995). Thus, citing Precedent 339 in its arguments, the State of Santa Catarina filed extraordinary appeals to the STF against lower court decisions that guaranteed to the servants (i) the right to the permanent increase, or (ii) equal salary adjustment to those servants receiving such an increase due to past commissioned positions and those who are currently in equivalent positions.
@@
To continue, one observes a significant rise in citations between 2003 and 2010, made of three peaks. They are all linked to specific events, represented by the presence of the words 10.192, emenda consitucional (constitutional amendment) or revisão geral anual (annual general review). Regarding the first one, a manual inspection revealed they refer to the application of Precedent 339 in the context of Law 10.192/2001, a law that increased the salaries of non-public servants (10,87%) and which was used by public servants to require a similar increase.
@@
As far as the 2005 peak is concerned, it must be traced back to 1993. Indeed, that was the year the Laws 8.622/1993 (already mentioned in Fig. 16) and 8.627/1993 came into effect, adjusting the salaries of higher ranks in the armed forces by 28.86%. STF Justices then recognized the same right to lower-ranking military personnel, understanding it was a case of general revision of salaries, therefore not applying Precedent 339 but rather Article 37, Section X of the Constitution. The Union filed extraordinary appeals [^51] Such as AI 551.744 https://portal.stf.jus.br/processos/detalhe.asp?incidente=2308620 (RE), and one of the arguments was that the Constitution allowed for different adjustments for different ranks (Article 142), but the Justices upheld their decision based on Article 37. Article 37 was amended by Constitutional Amendment 19/1998525252Constitutional Amendment 19/1998 https://www.planalto.gov.br/ccivil_03/constituicao/Emendas/Emc/emc19.htm#art3, which established the necessity of a specific law for any form of salary setting or change, thus exempting the requirement for adjustments to consistently follow the same percentage. In light of this, the Federal Government requested, and the STF accepted, that the 28.86% adjustment should be valid only until June 1998, when the Constitutional Amendment came into effect.
@@
We now turn to the last peak. In addition to the need for a specific law for any form of setting or change in the salary of public servants, Constitutional Amendment 19/1998 also established that the general revision of salaries must occur annually and based on annual laws proposed by the Head of the Executive branch (President of the Republic or state governors). A manual inspection of documents between 2006 and 2008 found extraordinary appeals [^53] Such as RE 524.145 https://portal.stf.jus.br/processos/detalhe.asp?incidente=2461438 requesting compensation for public servants due to omissions in issuing such a law. The STF denied these appeals based on Precedent 339. After many years of judgments on this matter, the STF stated in 2019 that not submitting a law’s proposition regarding the annual revision of public servants’ salaries does not create a subjective right to compensation; however, the Executive branch must provide justification for why the revision was not proposed (STF Tema 19545454Topic 19 https://portal.stf.jus.br/jurisprudenciaRepercussao/tema.asp?num=19[58].
@@
It has also created, due to the persistence of this resistance, an influx of cases to the higher courts, not only to STF but also to STJ (Superior Tribunal de Justiça), corresponding to the hypothesis New avenue to the court. As the controversy around this issue ensues, new cases continue to be brought before the STF, as new attempts to adjust the salary of public servants occur in the absence of formal laws. The contradiction of this never-ending administrative practice to the content of the BP, in itself, creates a clear avenue for the issue to be brought up time and time again before the court. In this case, however, it is interesting to note that this new avenue is not a consequence of lawyers trying to bend formal restrictions or take advantage of indirect legal connections to gain access to the court. Rather, this avenue is a result of direct confrontation between persevering administrative practices and what has been determined as the correct legal interpretation of the Law by the STF.